Shift Summary for owl shift, 28/aug/99

AG (ag@fnal.gov)
Sat, 28 Aug 1999 07:56:10 -0400 (EDT)

Shift Summary for owl shift, 28/aug/99

On Shift: BobT, AG

Hours of beam available at all intensities: 7h 30min
Hours of beam available at nominal intensity: 7h 30min
Hours of E832 data (special runs not included): 5h 24min
Hours of special runs (laser scans, muon runs...): 0h 0min

Accounting of lost time:

Debugging channel 450 gain drift.
Changing tapes.

========================================================================

Runs ended this shift:

Run 14351 - 18.9M - BEAM_832
Run 14353 - 2.6M - BEAM_832
Run 14354 is in progress.

Tapes brought to FCC this shift:
(including tapes for online split)

QKC336,QKC337,QKC338,
QKK336,QKC337,QKC338,
QKS336,QKS337,QKS338,
QKB117

========================================================================

Accesses during shift:

None.

========================================================================

Overview of shift:

Stable beam and smooth e8332 data taking, except for ~3 hours at the
beginning of the shift, when we tried to debug channel 450 gain drift.

========================================================================

Description of Shift:

run____spill______time_______comment

------------------------------------------------------------------------

14351____________0000______Shift starts with Val insisting on
___________________________paging CSI expert: CsI channel 450 has
___________________________gain drift of -15%.

_________________0010______Theo pages Sydney, who arrives shortly to
___________________________start nearly 3 hours adventure...

143351___________0101______We stop the run in order to tweak HV on channel
___________________________450.

14352________0120-0300_____Sydney,BobT and Ed (over the phone) work on
___________________________channel 450. Details can be found in ktev
___________________________logbook XXI, pp.100-103. Shortly, the result is
___________________________that (i) gain drift is still present and HV
___________________________change does not help, (ii) HV was changed from
___________________________-1546V to -1700V, (iii) dbase constants are
___________________________updated for this channel, (iv) problem is
___________________________likely a blown dynode on the HV divider, which
___________________________clearly should be swapped. But we decide to
___________________________wait until there is a downtime or the channel
___________________________becomes really erratic.

14353__________0301________Start new run.

14353__________0358________Tapes are full.

14354__________0415________Start new run.

14354____6_____0423________CsI channel 450 gain jumps by -6%.

________120____0657________CsI channel 450 is still stable at -6% wrt
___________________________dbase.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

o When starting new run, do dart fresh start to fix DC monitoring.

o Again, some of the online plots will look weird with the new trigger
configuration. They are marked in the plot book with sticky-notes.